OpenStudy (anonymous):

What is the graph of the rational function? y = 2x + 4 / x + 1

OpenStudy (anonymous):

im not sure whats going on but im not seeming to get a right answer for any of my graph questions

OpenStudy (anonymous):

y=(2x+4)/(x+1) = (2x+2+2)/(x+1) = 2 + { 2/(x+1)} So, y=2+{2/(x+1)}

OpenStudy (anonymous):

First try to draw graph of y=1/(x+1)

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

im not sure how to draw it thats why i was using my calculator

OpenStudy (anonymous):

@RitiBahl100 can you help me?

OpenStudy (anonymous):

@UnkleRhaukus can you help me?

OpenStudy (unklerhaukus):

do you know how to find the asymptotes ?